OpenStudy (anonymous):

√(2)*cos^2(x)+sin(x)=0 find solution\find x

OpenStudy (anonymous):

\[\sqrt{2}\times \cos x \times \cos x +\sin x =0\] please help

OpenStudy (cwrw238):

use the identity cos^2 x = 1 - sin^2 x to get a quadratic equation in sin x

OpenStudy (cwrw238):

follow?

OpenStudy (cwrw238):

plug in (1 - sin^2 x) for cos^2 x

OpenStudy (anonymous):

hey, didn't noticed you answered. thanks but emm, the quad equation that came out is 2t^2 - t -2 = 0 no concrete solutions to that..

OpenStudy (cwrw238):

√2 - √2 sin^2x + sin x = 0 √2 sin^2 x - sin x - √2 = 0

OpenStudy (anonymous):

yea, thats nice, but how do I continue...? hehe

OpenStudy (cwrw238):

if you solve this using the quadratic formula you get sin x = -1 / √2 or √2 as the sine of an angle must be between -1 and 1 the only possibility is - 1 / √2

OpenStudy (cwrw238):

so now you can find the possible values of x

OpenStudy (anonymous):

yes!, got it right... -45+360K \ 225+360K thanks alot dude :)

OpenStudy (cwrw238):

yw